OpenStudy (anonymous):

A little more help, anyone? This one is just like the last question I asked :( What is the solution set of |x – 6| = 10?

OpenStudy (jamesj):

Geometrically, this is asking "what are the points on the number line, x, that are exactly a distance 10 from the point 6" So draw a number line and figure it out. Tell me what you get then I'll show you the algebraic way to get the answer.

jimthompson5910 (jim_thompson5910):

or you can use this rule if |x| = k for some positive number k, then x = k or x = -k so... |x-6| = 10 turns into x-6 = 10 or x-6 = -10 solve each for x

OpenStudy (anonymous):

x=16 or x=-4

jimthompson5910 (jim_thompson5910):

good on both

OpenStudy (anonymous):

Awesome!